OpenStudy (anonymous):

]Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (cwrw238):

use the general form y - y1 = m(x - x1) m = slope which will be 3 (because its 3x in your equation) and (x1,y1) = (3, -2)

OpenStudy (cwrw238):

if you plug these values in and rearrange you'll get the correct option

OpenStudy (anonymous):

so i didvide 3 on both sides

OpenStudy (anonymous):

divide*

OpenStudy (cwrw238):

no you use the general equation above and replace m by 3 , x1 by 3 and y1 by -2 then rearrange to get y = whatever

OpenStudy (cwrw238):

i'll start you off: y - (-2) = 3(x - 3) y + 2 = 3x - 9 can you continue/

OpenStudy (anonymous):

ok now i get it yess i can contiune

OpenStudy (cwrw238):

good